the sum of john's age and Mary's age is 36 years. John is 4 years younger than Mary. how old are john and Mary?
Answer by Math_Boss(45)

J+M=36
J=(M-4)
Substitute J=(M-4)
(M-4)+M=36
M-4+M=36
2M=36+4
2M=40
M=20
J=20-4=16
Ans:Mary is 20, John is 16